OpenStudy (anonymous):

Convert: 35 mi/hr to km/min

OpenStudy (anonymous):

0.0268224 kilometers / minute So... 0.0268224*35 =0.938784km/min.

OpenStudy (mathstudent55):

\(35 \dfrac{mi}{hr} \times \dfrac{5280~ft}{mi} \times \dfrac{12 ~in.}{ft} \times \dfrac{2.54~cm}{in.}\times \dfrac{m}{100~cm} \times \dfrac{km}{1000~m} \times \dfrac{hr}{60~min}\) \(=0.94 \dfrac{km}{min}\)